# Provable Sample-Efficient Transfer Learning Conditional Diffusion Models via Representation Learning

### v1

@@ 3.1 Tackling Lipschitz Continuity under Weaker Assumptions
Notice that we only impose smoothness assumption on the original data distribution $p(\cdot|y)$, instead of the entire trajectory $p_{t}(\cdot|y)$ in forward process. This is substantially weaker than the Lipschitzness assumption required in Chen et al. [2023b, 2022b]; Yuan et al. [2024]; Yang et al. [2024]. However, Lipschitzness of loss function $\ell$ and class $\mathcal{F}$ is a crucial hypothesis in theoretical analysis of transfer learning [Tripuraneni et al., 2020; Chua et al., 2021]. The intuition is that without Lipschitz continuity of the score network $f$, it is generally impossible to characterize the error from an imperfect representation map $h$. Hence it is inevitable to show the smoothness of $p_{t}(\cdot|y)$ to some extent.
Fortunately, even with assumptions merely on the initial data distribution, we are still able to prove smoothness of the forward process in any bounded region, as shown in the following lemma. The proof can be found in Appendix A.1.
@@
Under Assumption 3.1, 3.2, 3.3, for any $w\in[0,1]^{d_{y}}$, denote the conditional score of forward process $\nabla_{x}\log p_{t}(x;w)$ by $f_{*}(x,w,t)$. There exist constants $C_{X},C_{X}^{\prime}$, such that for any $t\in[0,T]$, the function $f_{*}(x,w,t)$ is $(C_{X}+C_{X}^{\prime}\|x\|^{2})$-Lipschitz in $x$, $(C_{X}+C_{X}^{\prime}\|x\|)$-Lipschitz in $w$.
@@ 3.2 Results under Task Diversity: Sample-Efficient Transfer Learning
In the literature of transfer learning, task diversity is an important assumption that connects target tasks with source tasks [Tripuraneni et al., 2020; Du et al., 2020; Chua et al., 2021]. In the context of conditional diffusion models, we state the formal definition as follows.
@@
Here $L^{\mathbb{P}}$ is defined in (2.4). This notion of diversity ensures that the representation error on the target task caused by $\widehat{h}$ can be controlled by the error on the source tasks, thereby establishing certain relationships in between. More detailed discussions are deferred to Appendix A.5.
@@
$\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}}\mathbb{E}_{(x,y)\sim\mathbb{P}^{0}}[\ell^{\mathbb{P}^{0}}(x,y,s_{\widehat{f}^{0},\widehat{h}})]$ $\lesssim\frac{1}{\nu}\inf_{h\in\mathcal{H}}\frac{1}{K}\sum_{k=1}^{K}\inf_{f\in\mathcal{F}}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}[\ell^{\mathbb{P}^{k}}(x,y,s_{f,h})]+\Delta$ (3.4)
@@
The formal statements and proofs are provided in Appendix A.2.
@@
Comparison with Existing Transfer Learning Theory
Similar generalization bound of supervised transfer learning has been established under $(\nu,\Delta)$ diversity assumption. Tripuraneni et al. [2020] proves $\widetilde{\mathcal{O}}\left(\sqrt{\frac{K\log\mathcal{N}_{\mathcal{F}}+\log\mathcal{N}_{\mathcal{H}}}{nK}}+\sqrt{\frac{\log\mathcal{N}_{\mathcal{F}}}{m}}\right)$ in realizable case. Watkins et al. [2023] improves the rate to $\widetilde{\mathcal{O}}\left(\varepsilon_{\text{approx}}+\frac{K\log\mathcal{N}_{\mathcal{F}}+\log\mathcal{N}_{\mathcal{H}}}{nK}+\frac{\log\mathcal{N}_{\mathcal{F}}}{m}\right)$, by additionally assuming smoothness of loss function and applying local Rademacher complexity techniques. The difference in our analysis lies in the intricacy of time-dependent score matching loss, where the Lipschitzness and (or) smoothness need to be re-verified. Despite these technical difficulties, we are able to prove the same generalization bound as in supervised transfer learning.
@@ 3.3 Results without Task Diversity: Meta-Learning Perspective
[…]
@@
$\mathbb{E}_{\mathbb{P}^{0}\sim\mathbb{P}_{\text{meta}}}\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}\sim\mathbb{P}^{0}}\mathbb{E}_{(x,y)\sim\mathbb{P}^{0}}[\ell^{\mathbb{P}}(x,y,s_{\widehat{f}^{0},\widehat{h}})]$ (3.9)
@@
The formal statements and proofs are provided in Appendix A.3.
@@
Comparison with Existing Meta-learning Theory
The state of the art sample complexity bound in meta-learning setting is $\widetilde{\mathcal{O}}\left(\widetilde{\varepsilon}_{\text{approx}}+\sqrt{\frac{\log\mathcal{N}_{\mathcal{F}}}{m}}+\sqrt{\frac{\log\mathcal{N}_{\mathcal{H}}}{K}}\right)$ by assuming $m=n$ [Maurer et al., 2016], where the dependence on $\Omega(\sqrt{1/K})$ term cannot be improved. However, due to the smoothness of quadratic loss function, we are able to prove an even faster rate for score matching objective by leveraging local Rademacher complexity theory.
@@ 4.1 Score Neural Network Approximation
The following theorem provides a guarantee for the ability of deep ReLU neural networks to approximate score and representation. The proof is provided in Appendix B.1.
@@
[…]
@@
Universal approximation of deep ReLU neural networks in a bounded region has been widely studied [Yarotsky, 2017; Schmidt-Hieber, 2020]. However, we have to deal with an unbounded domain here, hence more refined analysis is required, e.g. truncation arguments.
In addition, traditional approximation theories typically cannot provide Lipschitz continuity guarantees, which is crucial in transfer learning analysis. Following the constructions in Chen et al. [2023b], the Lipschitzness restriction doesn’t compromise the approximation ability of neural networks, while ensuring validity of the generalization analysis in Section 3.
@@ 4.2 Distribution Estimation Error Bound
Given the approximation and generalization results, we are in the position of bounding the distribution estimation error of our transfer (meta) learning procedures. The formal statements and proofs can be found in Appendix B.2.
@@
Fu et al. [2024] studies conditional diffusion model for sub-gaussian distributions with $\beta$-Hölder density. Since the Lipschitzness of score is analogous to the requirement of twice differentiability of density [Wibisono et al., 2024], it is reasonable to let $\beta=2$ for a fair comparison. In this case, the TV distance is bounded by $\widetilde{\mathcal{O}}(m^{-\frac{1}{2(d_{x}+D_{y}+2)}})$ with sample size $m$ according to Fu et al. [2024], which is worse than our naive bound $\widetilde{\mathcal{O}}(m^{-\frac{1}{d_{x}+D_{y}+9}})$ due to the inefficiency of score approximation. We are also aware of another work [Jiao et al., 2024] that assumes Lipschitz density and score, obtaining a rate of $\widetilde{\mathcal{O}}(m^{-\frac{1}{2(d_{x}+3)(d_{x}+D_{y}+3)}})$.
Relation to Yang et al. [2024]
@@
Unlike our setup, Yang et al. [2024] considers transfer learning unconditional diffusion models with only one source task, i.e., $D_{y}=d_{y}=0,K=1$. The unconditional distribution is assumed to be supported in a low-dimensional linear subspace, where the source task and the target task have the same latent variable distribution. Hence, only a linear encoder is trained for fine-tuning instead of the full score network. In this case, Yang et al. [2024] is able to bound the TV distance by $\widetilde{\mathcal{O}}(m^{-\frac{1}{4}}+n^{-\frac{1-\alpha(n)}{d_{x}+5}})$, escaping the curse of dimensionality for target task. However, the assumption on shared latent variable distribution is stringent and we believe our analysis methods can be extended to this setting as well.
5 Applications
@@
5.1 Amortized Variational Inference
@@
$p_{\psi}(y|x)=\psi(y)\exp(\langle x,h_{*}(y)\rangle-A_{\psi}(x)),$ (5.1)
where $\psi\in\Psi$ is non-negative and supported on $[0,1]^{D}$ and $h_{*}(y)\in[0,1]^{d}$. Note that we also have $d_{x}=d$ in this case. The prior distribution of variable $x$ is denoted as $p_{\phi}$ for some $\phi\in\Phi$. Let $\theta=(\psi,\phi)$ and we aim to sample from the posterior distribution of $p_{\theta}(x|y)\propto p_{\phi}(x)p_{\psi}(y|x)\propto p_{\phi}(x)\exp(\langle x,h_{*}(y)\rangle-A_{\psi}(x))$. Due to the special structure, the posterior $p_{\theta}(x|y)$ only depends on the low-dimensional feature $h_{*}(y)$, shared across all $\theta\in\Theta:=\Psi\times\Phi$. This formulation encompasses various applications including independent component analysis [Comon, 1994], inverse problem [Song et al., 2021; Ajay et al., 2022] and variational Bayesian inference [Kingma, 2013].
Consider source tasks consisting of $\theta^{1},\cdots,\theta^{K}\in\Theta$, and for each $\theta^{k}$ we have $n$ i.i.d. samples $\{(x_{i}^{k},y_{i}^{k})\}_{i=1}^{n}$. For the target task $\theta^{0}$, we only have $m$ samples $\{(x_{i}^{0},y_{i}^{0})\}_{i=1}^{m}$. We conduct our transfer learning procedures to train a conditional diffusion models ^$\widehat{\mathbb{P}}_{\theta^{0}}(\cdot|y)$. For theoretical analysis, we further impose some assumptions on the probabilistic model as follows.
Assumption 5.1.
@@
Theorem 5.1.
Suppose Assumption 5.1 holds. Then under meta-learning setting, we have with probability no less than $1-\delta$,
$\mathbb{E}_{\theta^{0}}\mathbb{E}_{\{(x_{i}^{0},y_{i}^{0})\}_{i=1}^{m}}\mathbb{E}_{y\sim\mathbb{P}_{\theta^{0}}}[\mathrm{TV}(\widehat{\mathbb{P}}_{\theta^{0}}(\cdot|y),\mathbb{P}_{\theta^{0}}(\cdot|y))]\lesssim\frac{\log^{\frac{5}{2}}(nK/\delta)\log^{3}(m\wedge n)}{(m\wedge n)^{\frac{1}{2d+9}}}+\frac{\log^{2}(nK/\delta)}{K^{\frac{1}{D+2}}}.$ (5.2)
@@
$\mathbb{E}_{\{(x_{i}^{0},y_{i}^{0})\}_{i=1}^{m}}\mathbb{E}_{y\sim\mathbb{P}_{\theta^{0}}}[\mathrm{TV}(\widehat{\mathbb{P}}_{\theta^{0}}(\cdot|y),\mathbb{P}_{\theta^{0}}(\cdot|y))]\lesssim\frac{\log^{\frac{5}{2}}(nK/\delta)\log^{3}((m/\nu)\wedge n)}{\nu^{\frac{1}{2}}((m/\nu)\wedge n)^{\frac{1}{2d+9}}}+\frac{\log^{2}(nK/\delta)}{\nu^{\frac{1}{2}}(nK)^{\frac{1}{D+2}}}+\sqrt{\Delta}.$ (5.3)
The proof is deferred to Appendix C.1. We show that under mild assumptions, transfer (meta) learning diffusion models can improve the sample efficiency for target task in the context of amortized variational inference. This error bound can be further extended to establish guarantees for statistical inference such as moment prediction, uncertainty assessment, etc.
5.2 Behavior Cloning via Meta-Diffusion Policy
Although originally developed for image generation tasks, diffusion models have recently been extended to reinforcement learning (RL) [Janner et al., 2022; Chi et al., 2023; Wang et al., 2022], enabling the modeling of complex distributions of dynamics and policies. In the context of meta-RL, some works have further utilized diffusion models for planning and synthesis tasks [Ni et al., 2023; He et al., 2023]. In this application, we focus on a popular framework of behavior cloning, diffusion policy [Chi et al., 2023], which uses conditional diffusion models to learn multi-modal expert policies in high-dimensional state spaces. In such settings, the state often corresponds to visual observations of the robot’s surroundings, such as high resolution images, and thus typically share a low-dimensional underlying representation.

### v2

@@ 3.1 Tackling Lipschitz Continuity under Weaker Assumptions
Notice that we only impose smoothness assumption on the original data distribution $p(\cdot|y)$, instead of the entire trajectory $p_{t}(\cdot|y)$ in forward process. This is substantially weaker than the Lipschitzness assumption required in Chen et al. (2023b, 2022b); Yuan et al. (2024); Yang et al. (2024). However, Lipschitzness of loss function $\ell$ and class $\mathcal{F}$ is a crucial hypothesis in theoretical analysis of transfer learning (Tripuraneni et al., 2020; Chua et al., 2021). The intuition is that without Lipschitz continuity of the score network $f$, it is generally impossible to characterize the error from an imperfect representation map $h$. Hence it is inevitable to show the smoothness of $p_{t}(\cdot|y)$ to some extent.
Fortunately, even with assumptions merely on the initial data distribution, we are still able to prove smoothness of the forward process in any bounded region, as shown in the following lemma. The proof can be found in Appendix B.1.
@@
Under Assumption 3.1, 3.2, 3.3, for any $w\in[0,1]^{d_{y}}$, denote the conditional score of forward process $\nabla_{x}\log p_{t}(x;w)$ by $f_{*}(x,w,t)$. There exist constants $C_{X},C_{X}^{\prime}$, such that for any $R>0$, the function $f_{*}(x,w,t)$ is $(C_{X}+C_{X}^{\prime}R^{2})$-Lipschitz in $x$, $(C_{X}+C_{X}^{\prime}R)$-Lipschitz in $w$, in the domain $\mathcal{B}_{R}\times[0,1]^{d_{y}}\times[0,T]$. Here $\mathcal{B}_{R}$ denotes the ball with radius $R$ centered at the origin.
@@ 3.2 Results under Task Diversity: Sample-Efficient Transfer Learning
In the literature of transfer learning, task diversity is an important assumption that connects target tasks with source tasks (Tripuraneni et al., 2020; Du et al., 2020; Chua et al., 2021). In the context of conditional diffusion models, we state the formal definition as follows.
@@
Here $L^{\mathbb{P}}$ is defined in (2.4). This notion of diversity ensures that the representation error on the target task caused by $\widehat{h}$ can be controlled by the error on the source tasks, thereby establishing certain relationships in between. More detailed discussions are deferred to Appendix B.5.
@@
[…]
@@
The formal statements and proofs are provided in Appendix B.2.
@@ 3.3 Results without Task Diversity: Meta-Learning Perspective
Under Assumption 3.1, 3.2, 3.3, there exists constant $C_{P}$ such that for $\{\mathbb{P}^{k}\}_{k=1}^{K}\overset{\textit{i.i.d.}}{\sim}\mathbb{P}_{\text{meta}}$, with probability no less than $1-\delta$,
@@
[…]
@@
The formal statements and proofs are provided in Appendix B.3.
@@ 4.1 Score Neural Network Approximation
The following theorem provides a guarantee for the ability of deep ReLU neural networks to approximate score and representation. The proof is provided in Appendix C.1.
@@
$\displaystyle\inf_{h\in\mathcal{H}}\frac{1}{K}\sum_{k=1}^{K}\inf_{f\in\mathcal{F}}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}[\ell^{\mathbb{P}^{k}}(x,y,s_{f,h})]=\mathcal{O}\left(\log^{2}(nK/(\varepsilon\delta))\varepsilon^{2}\right),\text{ (transfer learning)}$ (4.1)
$\displaystyle\inf_{h\in\mathcal{H}}\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\inf_{f\in\mathcal{F}}\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{f,h})]=\mathcal{O}\left(\log^{2}(nK/(\varepsilon\delta))\varepsilon^{2}\right),\text{ (meta-learning)}$ (4.2)
@@
Universal approximation of deep ReLU neural networks in a bounded region has been widely studied (Yarotsky, 2017; Schmidt-Hieber, 2020). However, we have to deal with an unbounded domain here, hence more refined analysis is required, e.g. truncation arguments.
In addition, traditional approximation theories typically cannot provide Lipschitz continuity guarantees, which is crucial in transfer learning analysis. Following the constructions in Chen et al. (2023b), the Lipschitzness restriction doesn’t compromise the approximation ability of neural networks, while ensuring validity of the generalization analysis in Section 3.
@@ 4.2 Distribution Estimation Error Bound
Given the approximation and generalization results, we are in the position of bounding the distribution estimation error of our transfer (meta) learning procedures. The formal statements and proofs can be found in Appendix C.2.
@@
Fu et al. (2024) studies conditional diffusion model for sub-gaussian distributions with $\beta$-Hölder density. Since the Lipschitzness of score is analogous to the requirement of twice differentiability of density (Wibisono et al., 2024), it is reasonable to let $\beta=2$ for a fair comparison. In this case, the TV distance is bounded by $\widetilde{\mathcal{O}}(m^{-\frac{1}{2(d_{x}+D_{y}+2)}})$ with sample size $m$ according to Fu et al. (2024), which is worse than our naive bound $\widetilde{\mathcal{O}}(m^{-\frac{1}{d_{x}+D_{y}+9}})$ due to the inefficiency of score approximation. We are also aware of another work (Jiao et al., 2024) that assumes Lipschitz density and score, obtaining a rate of $\widetilde{\mathcal{O}}(m^{-\frac{1}{2(d_{x}+3)(d_{x}+D_{y}+3)}})$.
Relation to Yang et al. (2024)
@@
Unlike our setup, Yang et al. (2024) considers transfer learning unconditional diffusion models with only one source task, i.e., $D_{y}=d_{y}=0,K=1$. The unconditional distribution is assumed to be supported in a low-dimensional linear subspace, where the source task and the target task have the same latent variable distribution. Hence, only a linear encoder is trained for fine-tuning instead of the full score network. In this case, Yang et al. (2024) is able to bound the TV distance by $\widetilde{\mathcal{O}}(m^{-\frac{1}{4}}+n^{-\frac{1-\alpha(n)}{d_{x}+5}})$, escaping the curse of dimensionality for target task. However, the assumption on shared latent variable distribution is stringent and we believe our analysis methods can be extended to this setting as well.
5 Experiments
Our theoretical results can be readily applied in various real world settings. In Appendix A, we investigate amortized variational inference and behavior cloning utilizing our theories, providing statistical guarantees of practical applications of CDMs. In addition, we conduct experiments on both synthetic and real world data to numerically verify the sample efficiency of transfer learning.
$m$ 10 20 30 40 50 100
fine-tuning 14.47 3.68 2.45 1.82 1.9 0.91
train-from-scratch 21.99 10.61 5.71 2.38 1.77 1.04
Table 2: MSEs for $\beta_{0}=5.5$.
$m$ 10 20 30 40 50 100
fine-tuning 6.14 2.65 1.61 1.08 0.96 0.45
train-from-scratch 24.41 20.62 18.67 13.49 7.03 1.23
Table 3: MSEs for $\beta_{0}=15$.
Conditioned Diffusion
The first numerical example is the high-dimensional conditioned diffusion (Cui et al., 2016; Yu et al., 2023) arising from the following Langevin SDE
where $\beta>0$ and $w_{s}$ is a one-dimensional standard Brownian motion. The SDE (5.1) is discretized by the Euler-Maruyama scheme with a step size of $0.02$, which defines the prior distribution $p_{\beta}(x)$ for the (discretized) trajectory $x=\left(u_{0.02},u_{0.04},\ldots,u_{1.00}\right)^{\top}\in\mathbb{R}^{50}$. We consider a conditional Gaussian likelihood function, $p(y|x)=\mathcal{N}(Mx,I_{100}/4)$, where $M\in\mathbb{R}^{100\times 50}$ is a pre-defined projection matrix. Given a set of pre-selected $\{\beta_{k};1\leq k\leq K\}$ with $\beta_{k}=k$ and $K=10$, the $k$-th joint source distribution is given by $\mathbb{P}^{k}(x,y)=p_{\beta_{k}}(x)p(y|x)$. The target distribution $\mathbb{P}^{0}(x,y)$ is given by $\beta_{0}=5.5$ (in-domain) or $\beta_{0}=15$ (out-of-domain). More details are found in Appendix E.1.
We report the MSEs of the estimated posterior mean of $\mathbb{P}^{0}(x|y)$ on the test samples in Table 3 and 3. We see that across different values of $\beta$ and $m$, the fine-tuned models can provide significantly more accurate posterior mean estimations in most cases, suggesting the effectiveness of the representation map $\widehat{h}$ learned in the pre-training phase. Notably, as the number of fine-tuning samples $m$ increases, the performance gaps between fine-tuned models and train-from-scratch models get smaller, since more training samples yield more generalization benefits and thus less dependence on the pre-trained model. This is aligned with our theoretical results. We also notice a large variance among the results of different replicates, and attribute the slightly worse performance of fine-tuned models at $m=50,\beta_{0}=5.5$ to the potential randomness.
Image Restoration
For a real data experiment, we consider the image restoration task on MNIST. We have $K=9$ source tasks with $\mathbbm{P}^{k}(x,y)=p_{k}(x)p(y|x)$, where the prior $p_{k}(x)$ is the data distribution of the digit $k$ in the MNIST data set ($1\leq k\leq K$) and $p(y|x)=\mathcal{N}(x,I_{784}/4)$. The target task is $\mathbbm{P}^{0}(x,y)=p_{0}(x)p(y|x)$, where $p_{0}(x)$ is the data distribution of the digit 0. We use the full MNIST 1-9 data for pre-training which corresponds to $n=5000$. For the finetuning phase, we consider $m=10,20,30,40,50,100$ training samples and $100$ test samples from $\mathbbm{P}^{0}(x,y)$. More details can be found in Appendix E.2.
We report the MSEs between estimated posterior mean of $\mathbbm{P}^{0}(x,y)=p_{0}(x)p(y|x)$ and the ground truth sample $x$ on the test samples in Table 4. We see that for all fine-tuning sample sizes $m$, the results obtained by fine-tuning consistently outperform those obtained by training from scratch, indicating the benefits of transfer learning. Similarly to the experiment on conditioned diffusion, we also observe a reduced performance gap as $m$ increases.
$m$ 10 20 30 40 50 100
fine-tuning 0.3799 0.2846 0.2544 0.2406 0.2404 0.2268
train-from-scratch 0.4409 0.3180 0.2746 0.2551 0.2501 0.2344
Table 4: MSEs on the image restoration task.
6 Conclusion and Discussion
In this paper, we take the first step towards understanding the sample efficiency of transfer learning conditional diffusion models from the perspective of representation learning. We provide a generalization guarantee for transferring score matching in CDMs in different settings. We further establish an end-to-end distribution estimation error bound using deep neural networks. Two practical applications are investigated based on our theoretical results. We hope this work can motivate future theoretical study on the popular transfer learning paradigm in generative AIs.
Although this work provides the first statistical guarantee for transfer learning in CDMs, it has several limitations that we plan to address in future research. First, our theoretical results heavily rely on the task diversity notion introduced in Section 3.1, which can be challenging to verify in practice. While we provide some preliminary empirical evidence in Appendix B.5, a more fine-grained theoretical and empirical analysis will be essential for a deeper understanding of CDMs. Second, our analysis focuses on the ERM estimator, whereas in practice, fine-tuning typically starts from a pre-trained model and may employ techniques such as LoRA. Incorporating these settings would allow for an optimization-based perspective on the sample efficiency of transfer learning. Finally, in our current formulation, the sample efficiency gain arises from reducing the complexity associated with learning the conditional encoder. Consequently, our results primarily apply to CDMs in which the conditional encoder constitutes a substantial part of the overall model. Extending the theory to settings where this assumption does not hold is an important direction for future work.
Acknowledgements
This work was supported by National Natural Science Foundation of China (grant no. 12201014, grant no. 12292980 and grant no. 12292983). The research of Cheng Zhang was support in part by National Engineering Laboratory for Big Data Analysis and Applications, the Key Laboratory of Mathematics and Its Applications (LMAM) and the Key Laboratory of Mathematical Economics and Quantitative Finance (LMEQF) of Peking University. The authors are grateful for the computational resources provided by the High-performance Computing Platform of Peking University. The authors appreciate the anonymous NeurIPS reviewers for their constructive feedback.
Appendix A Applications
@@
A.1 Amortized Variational Inference
@@
$p_{\psi}(y|x)=\psi(y)\exp(\langle x,h_{*}(y)\rangle-A_{\psi}(x)),$ (A.1)
where $\psi\in\Psi$ is non-negative and supported on $[0,1]^{D}$ and $h_{*}(y)\in[0,1]^{d}$. Note that we also have $d_{x}=d$ in this case. The prior distribution of variable $x$ is denoted as $p_{\phi}$ for some $\phi\in\Phi$. Let $\theta=(\psi,\phi)$ and we aim to sample from the posterior distribution of $p_{\theta}(x|y)\propto p_{\phi}(x)p_{\psi}(y|x)\propto p_{\phi}(x)\exp(\langle x,h_{*}(y)\rangle-A_{\psi}(x))$. Due to the special structure, the posterior $p_{\theta}(x|y)$ only depends on the low-dimensional feature $h_{*}(y)$, shared across all $\theta\in\Theta:=\Psi\times\Phi$. This formulation encompasses various applications including independent component analysis [Comon, 1994], inverse problem [Song et al., 2021, Ajay et al., 2022] and variational Bayesian inference [Kingma, 2013].
Consider source tasks consisting of $\theta^{1},\cdots,\theta^{K}\in\Theta$, and for each $\theta^{k}$ we have $n$ i.i.d. samples $\{(x_{i}^{k},y_{i}^{k})\}_{i=1}^{n}$. For the target task $\theta^{0}$, we only have $m$ samples $\{(x_{i}^{0},y_{i}^{0})\}_{i=1}^{m}$. We conduct our transfer learning procedures to train a conditional diffusion models $\widehat{\mathbb{P}}_{\theta^{0}}(\cdot|y)$. For theoretical analysis, we further impose some assumptions on the probabilistic model as follows.
Assumption A.1.
@@
Theorem A.1.
Suppose Assumption A.1 holds. Then under meta-learning setting, we have with probability no less than $1-\delta$,
$\mathbb{E}_{\theta^{0}}\mathbb{E}_{\{(x_{i}^{0},y_{i}^{0})\}_{i=1}^{m}}\mathbb{E}_{y\sim\mathbb{P}_{\theta^{0}}}[\mathrm{TV}(\widehat{\mathbb{P}}_{\theta^{0}}(\cdot|y),\mathbb{P}_{\theta^{0}}(\cdot|y))]\lesssim\frac{\log^{\frac{5}{2}}(nK/\delta)\log^{3}(m\wedge n)}{(m\wedge n)^{\frac{1}{2d+9}}}+\frac{\log^{2}(nK/\delta)}{K^{\frac{1}{D+2}}}.$ (A.2)
@@
$\mathbb{E}_{\{(x_{i}^{0},y_{i}^{0})\}_{i=1}^{m}}\mathbb{E}_{y\sim\mathbb{P}_{\theta^{0}}}[\mathrm{TV}(\widehat{\mathbb{P}}_{\theta^{0}}(\cdot|y),\mathbb{P}_{\theta^{0}}(\cdot|y))]\lesssim\frac{\log^{\frac{5}{2}}(nK/\delta)\log^{3}((m/\nu)\wedge n)}{\nu^{\frac{1}{2}}((m/\nu)\wedge n)^{\frac{1}{2d+9}}}+\frac{\log^{2}(nK/\delta)}{\nu^{\frac{1}{2}}(nK)^{\frac{1}{D+2}}}+\sqrt{\Delta}.$ (A.3)
The proof is deferred to Appendix D.1. We show that under mild assumptions, transfer (meta) learning diffusion models can improve the sample efficiency for target task in the context of amortized variational inference. This error bound can be further extended to establish guarantees for statistical inference such as moment prediction, uncertainty assessment, etc.
A.2 Behavior Cloning via Meta-Diffusion Policy
Although originally developed for image generation tasks, diffusion models have recently been extended to reinforcement learning (RL) [Janner et al., 2022, Chi et al., 2023, Wang et al., 2022], enabling the modeling of complex distributions of dynamics and policies. In the context of meta-RL, some works have further utilized diffusion models for planning and synthesis tasks [Ni et al., 2023, He et al., 2023]. In this application, we focus on a popular framework of behavior cloning, diffusion policy [Chi et al., 2023], which uses conditional diffusion models to learn multi-modal expert policies in high-dimensional state spaces. In such settings, the state often corresponds to visual observations of the robot’s surroundings, such as high resolution images, and thus typically share a low-dimensional underlying representation.